# Auto-ICL: In-Context Learning without Human Supervision

### v2

@@ 1 Introduction
In the era of Large Language Models (LLMs), human-computer interaction has evolved towards natural language, offering unprecedented flexibility through in-context learning (Devlin et al., 2018; Radford et al., 2019; Brown et al., 2020; Wei et al., 2022a). Given an appropriate context, the performance of LLMs can be significantly enhanced. Vanilla in-context learning relies on human-provided contexts, such as labeled examples (Few-Shot (Brown et al., 2020) and Few-Shot-CoT (Wei et al., 2022b) methods), explicit instructions, or other guiding mechanisms that shape the model’s outputs. However, human-written contexts present two major limitations:
@@ 2 Method
where $r$ is available context resource. We categorize Auto-ICL on whether we can retrieve [^2] Retrieving only employs queries without labels from the dataset, which is different from utilizing labeled examples in the Few-Shot method. similar queries from the dataset: (1) Auto-ICL (retrieving): similar queries from the same dataset are available, represented as $Q=\{q_{1},\ldots,q_{n}\}$, where we use them to assist context generation (i.e. $r=Q$). (2) Auto-ICL (generating): only the current query $q$ is available, and context is generated using this query (i.e. $r=q$).
@@ 2.1 Demonstration
Figure 2: Left: A simple example to generate demonstration by the model. Right: Given queries and reasoning paths, the model generates a general instruction.
@@
Generating demonstrations based on query $q$. The most straightforward case involves using the question $q$ to generate a demonstration composed of a set of queries $\hat{Q}$ and their corresponding label $\hat{A}$, denoted as $q\rightarrow(\hat{Q},\hat{A})$. For better performance, we directly generate demonstration with reasoning path toward the answer (chain-of-thought), represented as $q\rightarrow(\hat{Q},\hat{CoT})$. A simple example is shown in Figure 2
Retrieving similar queries $Q$ from dataset contain query $q$. On the other hand, when we have access to a dataset containing queries similar to $q$, we can retrieve a subset $Q$ from this dataset. The model then generates answers with reasoning path, denoted as $Q\rightarrow(Q,\hat{CoT})$, which is exactly the Auto-CoT method (Zhang et al., 2022). Hence, Auto-CoT can be viewed as a specific case under our framework where demonstrations with reasoning processes are generated from the retrieved dataset.
@@ 2.2 Instruction generation
Generating instruction based on retrieved queries $Q$ For each query in the set $\hat{Q}=\{\hat{Q}_{1},\ldots,\hat{Q}_{n}\}$ or $Q=\{q_{1},\ldots,q_{n}\}$, we generate an reasoning path, resulting in a series of individual chian-of-thought, denoted as $\hat{CoT}=\{\hat{CoT}_{1},\ldots,\hat{CoT}_{n}\}$. Then we synthesize these individual paths into a single, general instruction, $(Q,\hat{CoT})\rightarrow\hat{i}$. The instruction assures each reasoning step in the answer is consistent and adaptable to all queries within the dataset. Since multiple instructions are consolidated into one general instruction, this method is referred to as N-to-1 instruction generation. The method is visualized in Figure 2, and an actual instruction generated by GPT-3.5-Turbo-0301 is shown in Table 2.
As instructions can be generated upon demonstrations, we also consider incorporating both the demonstration and instruction into a context in the experiment.
@@ 3 Experiment Setting
Dataset To evaluate our method, we conduct experiments across four categories of tasks: Theory of Mind (Sap et al., 2019; Le et al., 2019; Sap et al., 2022), symbolic reasoning (Roy and Roth, 2016; Ling et al., 2017; Cobbe et al., 2021), arithmetic (Wei et al., 2022b; Kojima et al., 2022), and other reasoning tasks (Srivastava et al., 2022; Brown et al., 2020). Details of each tasks are listed below:
@@
Instruction N-to-1 generation: $(Q,CoT)\rightarrow\hat{i}$
Input:
……Demonstrations with chain-of-thought ……
Output:
@@
Table 3: The table compares our method with examplar-only methods and Few-Shot-CoT method across various models on the AQUA dataset. Auto-ICL (generating) refers to "Demonstration + Instruction" method and Auto-ICL (retrieving) refers to "Instruction" method in the Table 1. We use 5-shot examples in the demonstration.
Methods GPT-4 GPT-4-Turbo GPT-3.5-Turbo-0301 Llama-3-70B Avg.
5-Shot-CoT 88.5 83.0 73.0 84.0 82.1
@@
Table 4: The table compares our method with with examplar-only methods and Few-Shot-CoT method across various models on the GSM8K dataset. Auto-ICL (generating) refers to "Demonstration + Instruction" method and Auto-ICL (retrieving) refers to "Instruction" method in the Table 1. We use 5-shot examples in the demonstration.
Baselines We compared our method against baseline methods:
@@
Model We evaluate our method on the GPT family (Brown et al., 2020), including advanced models such as GPT-4, GPT-4-Turbo (evaluated during 2024 June), and GPT-3.5-Turbo-0301, as well as earlier models Text-Ada-001, Text-Davinci-003, Text-Davinci-002, and Text-Davinci-001. Additionally, we assess our approach on Llama-3-70B, Llama-2-13B, and PaLM2 text-bison (Anil et al., 2023). We set temperature as 0 for all models and tasks.
@@ 4 Results
We tested demonstration-only, instruction-only, and demonstration-plue-instruction contexts on the GPT-3.5-Turbo-0301 model across various tasks and show the average accuracy of each task in Table 1 (full result is in Appendix A.1). In the retrieving case, the instruction-only context outperformed other contexts, which we denote as Auto-ICL (retrieving) from Table 4 to Table 6. The demonstration-plus-instruction context yielded better results in the generating case, which we denote as Auto-ICL (generating) from Table 4 to Table 6. We have following observations:
Generated context outperforms human annotated context. Table 1, Table 6, and Table 6 demonstrate that our method surpasses traditional methods, including Zero-CoT and Auto-CoT. It also outperforms traditional context generation methods using labeled data. Tables 4 and 4 show that our method outperforms the Few-Shot-CoT method and concurrent work that uses LLMs to generate demonstrations as context only.
@@
Zero-CoT 70.5 73.0 37.4 18.0 72.0 53.0 45.1
Auto-CoT 75.5 64.0 53.0 18.5 46.5 54.8 43.2
Auto-ICL (generating) 73.0 66.0 52.5 33.3 66.5 58.5 52.7
Auto-ICL (retrieving) 73.0 71.4 60.8 15.0 71.0 52.5 49.8
Zero-CoT 91.5 90.7 56.9 0.5 91.3 72.0 62.3
Auto-CoT 92.5 89.3 78.0 1.5 78.3 71.0 63.6
Auto-ICL (generating) 93.5 90.0 78.0 2.5 91.8 74.0 67.3
Auto-ICL (retrieving) 92.0 91.4 79.0 3.0 93.3 72.0 67.7
Table 6: The table compares our method with baseline methods across various models on the GSM8K dataset. Auto-ICL (generating) refers to "Demonstration + Instruction" method and Auto-ICL (retrieving) refers to "Instruction" method in the Table 1. We use 5-shot examples in the demonstration.
In retrieving methods, instruction approaches consistently outperform demonstration and demonstration-plus-instruction methods in Tbale 1. This highlights the effectiveness of instructions in fine-tuning the model’s performance and guiding its problem-solving capabilities. Since the source of information remains the same (i.e., retrieved data), the repetitive usage of the same information in different formats (i.e., demonstration-plus-instruction) does not necessarily enhance the model’s problem-solving ability.
In generating methods, demonstration-plus-instruction outperformes instruction-only and demonstration-only approaches in Tbale 1. Combining demonstration and instruction proved to be more effective in generating methods than the independent use of instructions or demonstrations. Given that the information from a single question, q, may be limited, it can be inferred that expressing limited information in different formats—such as demonstration-plus-instruction—can enrich and extend the information used to address question $q$.
Limitation on earlier and smaller models. In Table 7, we evaluate our methods on earlier OpenAI models and the smaller Llama2-13B model. However, these methods do not perform as well compared to their performance on advanced models. This is because our approach heavily relies on the generation capabilities of the model. While it excels with the latest advanced models, it can negatively impact the performance of less capable models compared to the Zero-CoT method.
@@
Zero-CoT 52.5 36.5 17.5 23.5 32.5
Few-Shot 13.0 11.5 5.5 3.5 8.4
Auto-CoT 59.3 52.0 13.6 19.5 36.1
@@
5 Related Work
Self-generation of context for problem-solving has been progressively developed through multiple studies. Kim et al. (2022) firstly uses LLMs to generate demonstration queries in classification tasks. However, their work is limited by providing the model with the label and generating the corresponding demonstration. Further, Honovich et al. (2022); Zhou et al. (2022) leverage demonstration pairs to generate task instructions. Recently, Li et al. (2023) introduced a step-wise context generation approach, self-prompting, primarily applied in Open-Domain Question Answering. They focused solely on the QA dataset, proposing a method limited to generating demonstrations. Shao et al. (2023) use human-written demonstrations to prompt the model to generate additional demonstrations, followed by a selection process. Contrary to these prior studies, our method feeds the model with only a query, enabling it to generate demonstrations and instructions. Our research further proposes a method to convert a generated demonstration into an instruction format. Thus, our method is more general, requiring only queries, and can shown to be successful in Theory of Mind and reasoning tasks.
In knowledge-intensive tasks necessitating domain or world knowledge, Yu et al. (2022) and Sun et al. (2022) are geared towards using self-generated context to aid the model in extracting the correct answer from a document. In contrast, our method steers the model towards the right reasoning path by providing demonstrations and instructions with various generation paths rather than merely supplying the document containing the answer.
@@ 6 Discussion
From task-specific to universal in-context learning Recently, significant research has been conducted on the design of prompts to enhance the performance of LLMs on specific tasks (Li et al., 2023; Hu et al., 2023; Wang et al., 2023; Sun et al., 2023). However, these task-specific methodologies often suffer from two main limitations: first, they require a degree of pre-existing knowledge about the task to design an effective prompt, and second, they are typically not transferable across different tasks, limiting their versatility. Therefore, instead of creating task-specific prompts, our method uses a single, broad strategy applicable to various scenarios. The model can process the context of the problem and generate relevant demonstrations or instructions, enabling it to tackle many tasks. This approach represents a shift from task-specific to universal ICL.
Considering resources available under an in-context setting Our research serves as an example in assessing available in-context resources for problem-solving while also proposing and comparing multiple ways to utilize these resources. This work focused on the impact of demonstrations and instructions, formats that have proven successful in ICL. However, other forms of ICL might also enhance LLM performance. Future research could delve into alternative formats such as rephrasing questions, enriching context with anecdotes or narratives, or applying different reasoning processes. Moreover, our research primarily accessed one form of resource - data retrieval from a dataset. There is room to consider other resources like real-time data from the web or integrating user feedback during the learning process. Each of these potential resources could provide additional context or information, further enhancing the problem-solving abilities of LLMs.
@@ 8 Limitations
In our research, we develop Auto-ICL, a novel framework that allows LLMs to autonomously generate useful context to address user queries. Nonetheless, the efficacy of our technique hinges on the model’s inherent capacity to comprehend the present query and generate a pertinent and beneficial context for enhanced problem-solving. As we shown in the Section 4, the capability is predominantly found in advanced models.

### v3

@@ 1 Introduction
In the era of Large Language Models (LLMs), human-computer interaction has evolved towards natural language, offering unprecedented flexibility through in-context learning (Radford et al., 2019; Brown et al., 2020; Wei et al., 2022a). Given an appropriate context, the performance of LLMs can be significantly enhanced. Vanilla in-context learning relies on human-provided contexts, such as labeled examples (Few-Shot (Brown et al., 2020) and Few-Shot-CoT (Wei et al., 2022b) methods), explicit instructions, or other guiding mechanisms that shape the model’s outputs. However, human-written contexts present two major limitations:
@@
Figure 2: Left: An example to generate demonstration by the model. Right: Given queries and reasoning paths, the model generates a general instruction. To generate instructions, demonstrations can from retrieved set or generated set.
2 Related Work
Self-generation of context for problem-solving has been progressively developed through multiple studies. Kim et al. (2022) firstly uses LLMs to generate demonstration queries in classification tasks. However, their work is limited by providing the model with the label and generating the corresponding demonstration. Further, Honovich et al. (2022); Zhou et al. (2022) leverage demonstration pairs to generate task instructions. Recently, Li et al. (2023) introduced a step-wise context generation approach, self-prompting, primarily applied in Open-Domain Question Answering. They focused solely on the QA dataset, proposing a method limited to generating demonstrations. Shao et al. (2023) use human-written demonstrations to prompt the model to generate additional demonstrations, followed by a selection process. Contrary to these prior studies, our method feeds the model with only a query, enabling it to generate demonstrations and instructions. Our research further proposes a method to convert a generated demonstration into an instruction format. Thus, our method is more general, requiring only queries, and can shown to be successful in Theory of Mind and reasoning tasks. In knowledge-intensive tasks necessitating domain or world knowledge, Yu et al. (2022) and Sun et al. (2022) are geared towards using self-generated context to aid the model in extracting the correct answer from a document. In contrast, our method steers the model towards the right reasoning path by providing demonstrations and instructions with various generation paths rather than merely supplying the document containing the answer.
@@ 3 Method
where $r$ is available context resource. We categorize Auto-ICL on whether we can retrieve [^2] Retrieving only employs queries without labels from the dataset, which is different from utilizing labeled examples in the Few-Shot method. queries from the dataset: (1) Auto-ICL (retrieving): A dataset containing $q$ is available in this case. We use a small subset represented as $Q=\{q_{1},\ldots,q_{n}\}$, to generate context (i.e. $r=Q$). Once a context is generated, it is applied to the whole dataset. (2) Auto-ICL (generating): only the current query $q$ is available, and context is generated using this query (i.e. $r=q$).
@@ 3.1 Demonstration
Figure 3: An illustration of how to use generated context for problem-solving in Auto-ICL (retrieving) and Auto-ICL (generating). We place the Few-Shot-CoT method in the upper-left corner for comparison. In Auto-ICL (retrieving), the instruction is applied to all queries in the same dataset. In Auto-ICL (generating), the examples and instructions are tailored from each query. The generation process of demonstrations and instructions are shown in Figure 2.
@@
Generating demonstrations based on query $q$. The most straightforward case involves using the question $q$ to generate a demonstration composed of a set of queries $\hat{Q}$ and their corresponding label $\hat{A}$, denoted as $q\rightarrow(\hat{Q},\hat{A})$. For better performance, we directly generate demonstration with reasoning path toward the answer (chain-of-thought), represented as $q\rightarrow(\hat{Q},\hat{CoT})$. A simple example is shown in Figure 2.
Retrieving queries $Q$ from the dataset containing query $q$. On the other hand, when we have a dataset containing query $q$, we can randomly retrieve a subset $Q$ from this dataset. The model then generates answers with reasoning path, denoted as $Q\rightarrow(Q,\hat{CoT})$, which is exactly the Auto-CoT method (Zhang et al., 2022). Hence, Auto-CoT can be viewed as a specific case under our framework where demonstrations with reasoning processes are generated from the retrieved dataset.
@@ 3.2 Instruction generation
Generating instruction based on queries. Start from demonstration generation, for each query in the randomly retrieved subset $Q$ or generated subset $\hat{Q}$, we generate a set of reasoning path, denoted as $\hat{CoT}$. Then we synthesize these individual paths into a single, general instruction, $(Q,\hat{CoT})\rightarrow\hat{i}$ or $(\hat{Q},\hat{CoT})\rightarrow\hat{i}$. The instruction assures each reasoning step in the answer is consistent and applicable to all queries within the dataset. Since multiple instructions are consolidated into one general instruction, this method is referred to as N-to-1 instruction generation. The method is visualized in Figure 2, and an actual instruction generated by GPT-3.5-turbo-0301 is shown in AppendixA.2.2. As instructions can be generated upon demonstrations, we also consider incorporating both the demonstration and instruction into a context in the experiment.
3.3 Auto-ICL (retrieving) and Auto-ICL (generating)
Since the context can be a combination of demonstrations, instructions, or each on its own, we summarize two forms under the condition of whether we have a dataset to retrieve queries based on the results from experiments:
Auto-ICL (retrieving): With retrieval queries, we suggest using instructions only. Here, demonstrations are retrieved from the dataset, answers are generated with Zero-CoT, and a single instruction is created. This instruction is then applied to all queries within the same dataset.
Auto-ICL (renerating): Without retrieval queries, we use both demonstrations and instructions. The process starts by generating demonstrations for the query $q$, followed by generating an instruction using these demonstrations.
After creating the context, we combine it with the query $q$ and input them into the model. An example is provided in Figure 3.
3.4 Comparison of computational efficiency
In Auto-ICL (retrieving), the generated context is general and applicable to queries within the same dataset. Consequently, the decoding cost is comparable to that of the Zero-CoT and Few-Shot-CoT methods. Additionally, the generated instruction is much shorter than the demonstrations used in Few-Shot-CoT cases as the number of shots increases, thereby reducing encoding costs.
In Auto-ICL (generating), we assume access only to the current query without similar query examples, reflecting real-world use cases where users may have only a single query. Here, the context is specifically generated for each query. Our method incurs additional decoding costs compared to the Zero-CoT method but results in higher accuracy.
Compared to APE for instruction generation (Zhou et al., 2022), which requires generating several instructions and iterating them to maximize a score function, our method is computationally efficient. APE also needs access to the log probabilities of models, which our method does not require.
In conclusion, Auto-ICL is a simple and cost-effective method that achieves higher accuracy in the retrieving case. Although there is an accuracy-cost trade-off in the generating case compared to Zero-CoT, Auto-ICL does not replies on other resources compared with other methods.
Figure 4: A comparison of instruction generated by different methods on AQUA dataset with GPT-3.5-turbo-0301.
@@ 4 Experiment Setting
Dataset: To evaluate our method, we conduct experiments across four categories of tasks: Theory of Mind (Sap et al., 2019; Le et al., 2019; Sap et al., 2022), symbolic reasoning (Roy and Roth, 2016; Ling et al., 2017; Cobbe et al., 2021), arithmetic (Wei et al., 2022b; Kojima et al., 2022), and other reasoning tasks (Srivastava et al., 2022; Brown et al., 2020). Details of each tasks are listed below:
@@
Model: We evaluate our method on the GPT family (Brown et al., 2020), including advanced models such as GPT-4, GPT-4-Turbo (evaluated during 2024 June), and GPT-3.5-Turbo-0301, as well as earlier models Text-Ada-001, Text-Davinci-003, Text-Davinci-002, and Text-Davinci-001. Additionally, we assess our approach on Llama-2-13B (Touvron et al., 2023), Llama-3-70B (AI@Meta, 2024), and PaLM2 text-bison (Anil et al., 2023). We set temperature as 0 for all models and tasks.
@@
5-Shot-CoT (Brown et al., 2020) 88.5 83.0 73.0 84.0 82.1
@@
Table 2: The table compares our method with examplar-only methods and Few-Shot-CoT method across various models on the AQUA dataset (top) and GSM8K dataset (bottom). Auto-ICL (generating) refers to "Demonstration + Instruction" method and Auto-ICL (retrieving) refers to "Instruction" method in the Table 1. We use 5-shot examples in the demonstration.
Baseline: We compared our method against baseline methods:
@@
7. DSPy frameworks (Khattab et al., 2023) enabling few-shot examples bootstrapping. We compared our method with the DSPy framework using the GSM8K dataset on GPT-3.5-turbo and GPT-4-turbo  [^3] As GPT-3.5-turbo-0301 was deprecated when we drew attention to this work during reviewing, we used the latest model instead. For the DSPy framework, we enabled the 5-Shot-CoT demonstration selection with 50 training examples and 200 development examples on GPT-3.5-turbo, 30 training examples, and 50 development examples on GPT-4-turbo.
@@
Table 3: Comparison of Auto-ICL methods with DSPy framework on the GSM8K dataset.
Method GPT-3.5-Turbo GPT-4-Turbo
DSPy (Khattab et al., 2023) 81.0 93.0
Auto-ICL (generating) 82.0 90.0
Auto-ICL (retrieving) 83.0 91.4
Zero-CoT (Brown et al., 2020) 70.5 73.0 37.4 18.0 72.0 53.0 45.1
Auto-CoT (Zhang et al., 2022) 75.5 64.0 53.0 18.5 46.5 54.8 43.2
Auto-ICL (generating) 73.0 66.0 52.5 33.3 66.5 58.5 52.7
Auto-ICL (retrieving) 73.0 71.4 60.8 15.0 71.0 52.5 49.8
Zero-CoT (Brown et al., 2020) 91.5 90.7 56.9 0.5 91.3 72.0 62.3
Auto-CoT (Zhang et al., 2022) 92.5 89.3 78.0 1.5 78.3 71.0 63.6
Auto-ICL (generating) 93.5 90.0 78.0 2.5 91.8 74.0 67.3
Auto-ICL (retrieving) 92.0 91.4 79.0 3.0 93.3 72.0 67.7
@@ 5 Results
We tested demonstration-only, instruction-only, and demonstration-plue-instruction contexts on the GPT-3.5-Turbo-0301 model across various tasks and show the average accuracy of each task in Table 1 (full result is in Appendix A.1). In the retrieving case, the instruction-only context outperformed other contexts, which we denote as Auto-ICL (retrieving) from Table 2. The demonstration-plus-instruction context yielded better results in the generating case, which we denote as Auto-ICL (generating) from Table 2.
5.1 Comparison with baseline methods
Generated context outperforms human annotated context. Table 1 and Table 4 demonstrate that our method surpasses traditional methods, including Zero-CoT and Auto-CoT. It also outperforms traditional context generation methods using labeled data. Table 2 shows that our method outperforms the Few-Shot-CoT method and concurrent work that uses LLMs to generate demonstrations as context only. We also compare our method with demonstration selection with DSPy framework in Table 3, our method demonstrates performance similar to the DSPy framework. Additionally, the DSPy framework consumes a significant amount of credits for iterating through demonstrations in the training set, and it relies on demonstrations with labels. In contrast, our method substantially reduces inference costs and does not depend on human-written labels.
@@ 5.1 Comparison with baseline methods
Zero-CoT (Brown et al., 2020) 52.5 36.5 17.5 23.5 32.5
Few-Shot (Brown et al., 2020) 13.0 11.5 5.5 3.5 8.4
Auto-CoT (Zhang et al., 2022) 59.3 52.0 13.6 19.5 36.1
@@
Instruction Interpretability: We present a comparison of generated instructions by APE, Instruction Induction, and Auto-ICL in Figure 4 In this example, the Instruction Induction method fails to generate an effective instruction, while APE relies on a specific example. Our method produces a more general and applicable instruction. In conclusion, our method’s explanations (instructions) on problems are more interpretable by human.
5.2 Are retrieved queries useful?
5.3 Instruction vs. Demonstration
In retrieving methods, instruction approaches consistently outperform demonstration and demonstration-plus-instruction methods in Table 1. This highlights the effectiveness of instructions in fine-tuning the model’s performance and guiding its problem-solving capabilities. Since the source of information remains the same (i.e., retrieved data), the repetitive usage of the same information in different formats (i.e., demonstration-plus-instruction) does not necessarily enhance the model’s problem-solving ability.
In generating methods, demonstration-plus-instruction outperformes instruction-only and demonstration-only approaches in Table 1. Combining demonstration and instruction proved to be more effective in generating methods than the independent use of instructions or demonstrations. Given that the information from a single question, q, may be limited, it can be inferred that expressing limited information in different formats—such as demonstration-plus-instruction—can enrich and extend the information used to address question $q$.
@@ 6 Discussion
From task-specific to universal in-context learning. Recently, significant research has been conducted on the design of prompts to enhance the performance of LLMs on specific tasks (Li et al., 2023; Hu et al., 2023; Wang et al., 2023; Sun et al., 2023). However, these task-specific methodologies often suffer from two main limitations: first, they require a degree of pre-existing knowledge about the task to design an effective prompt, and second, they are typically not transferable across different tasks, limiting their versatility. Therefore, instead of creating task-specific prompts, our method uses a single, broad strategy applicable to various scenarios. The model can process the context of the problem and generate relevant demonstrations or instructions, enabling it to tackle many tasks. This approach represents a shift from task-specific to universal ICL.
Considering resources available under an in-context setting. Our research serves as an example in assessing available in-context resources for problem-solving while also proposing and comparing multiple ways to utilize these resources. This work focused on the impact of demonstrations and instructions, formats that have proven successful in ICL. However, other forms of ICL might also enhance LLM performance. Future research could delve into alternative formats such as rephrasing questions, enriching context with anecdotes or narratives, or applying different reasoning processes. Moreover, our research primarily accessed one form of resource - data retrieval from a dataset. There is room to consider other resources like real-time data from the web or integrating user feedback during the learning process. Each of these potential resources could provide additional context or information, further enhancing the problem-solving abilities of LLMs.
@@ 8 Limitations
In Table 5, we evaluate our methods on earlier OpenAI models, including Text-Davinci-001, 002, and 003, as well as the smaller Llama2-13B model. However, these methods do not perform as well compared to their performance on advanced models. This discrepancy arises because our approach heavily relies on the generation capabilities of the model. In addition, compared to GPT-3.5-Turbo and GPT-4, the limitations of the Davinci series, such as a lack of instruction comprehension and sensitivity to prompt variations, can negatively impact performance. As a result, Auto-ICL underperforms with less capable models compared to the Zero-CoT method.